OpenStudy (anonymous):

What is the equation of curve of intersection of sphere x^2+y^2+z^2=a^2 and x+z=a

OpenStudy (dumbcow):

use substitution \[x^{2}+y^{2}+z^{2} = (x+z)^{2}\] \[\rightarrow y^{2} = 2xz\] z = a - x \[y = \pm \sqrt{2x(a-x)}\]